128.6°F to Celsius scale

1
Expert's answer
2022-04-11T12:46:44-0400
"\rm C=\frac{5}{9}(F-32)"

"\rm C=\frac{5}{9}(128.6-32)=53.7\:\rm {^\circ C}"